OpenStudy (anonymous):

Find the number of real solutions of the quadratic equation with: a = 6, b = 3, c = 2. A. no real solutions B. 1 real solution C. 2 real solutions D. 3 real solutions

OpenStudy (anonymous):

its A

jimthompson5910 (jim_thompson5910):

you are correct, hopefully you see how to get that?

OpenStudy (anonymous):

Quadratic equations have no solution when the quadratic formula gives non-real answers only. This is due to the \[\sqrt{b^{2}-4ac}\] term. If there's a negative number under that root, the solutions are unreal. Can you use that to explain when quadratics have only one root (solution)?

OpenStudy (anonymous):

cause they dont give you information or detail on how to do it @jim_thompson5910

jimthompson5910 (jim_thompson5910):

no you can use the given info to determine how many real solutions you'll get

OpenStudy (anonymous):

yeah but i mean like they only give you answers

jimthompson5910 (jim_thompson5910):

you use the discriminant formula D = b^2 - 4ac D = 3^2 - 4(6)(2) ... plug in a = 6, b = 3, c = 2 D = 9 - 48 D = -39 Because D < 0, this means that you'll have no real solutions (and 2 complex solutions instead)

jimthompson5910 (jim_thompson5910):

If D = 0, then you have exactly one real solution (a rational solution) If D > 0, then you'll have 2 distinct real solutions